how do you solve -2n + 14 = 34

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Chriss,

-2n+14=34

2n=14-34=-20

n=-10

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Chris !!!

 


Regards,

Askiitians Experts

CHETAN MANDAYAM NAYAKAR